# Yuehao Bai

### v7

@@ 6 Recommendations for Empirical Practice
We conclude with some recommendations for empirical practice based on our theoretical results. Overall, our findings highlight the general benefit of fine stratification for designing efficient experiments: finely stratified experiments “automatically” perform fully-efficient covariate adjustment for a large class of interesting parameters. This finding generalizes similar observations made by bai2022inference, bai2022optimality and cytrynbaum2023designing for the special case of estimating the ATE.
Our simulation evidence suggests, however, that finely stratified experiments may produce less precise estimates than (correctly specified) covariate adjustment when the the dimension of $X_{i}$ is large relative to the sample size. For this reason, we recommend that practitioners construct their blocks using a subset of the baseline covariates that they believe have the highest explanatory power in terms of the nonparametric $R^{2}$ in (19); the pre-treatment measure of the outcomes of interest, for example, is typically believed to be one such covariate (see, in particular, bruhn2009pursuit). The experimental data can then be analyzed efficiently using an unadjusted method-of-moments estimator.
If one wishes to perform covariate adjustment with additional covariates beyond those used for blocking, then this can be done ex post. As discussed in Remark 4.2, the scope for improvement from covariate adjustment is limited by the nonparametric $R^{2}$ from the regression of the moment functions on the additional covariates, conditional on the ones used for matching; if one has already matched on the covariates with the highest explanatory power, then the potential gain in efficiency from adjusting for these additional covariates may be limited. We further caution that care must be taken to ensure that the adjustment is performed in such a way that it guarantees a gain in efficiency: see bai2024covariate and cytrynbaum2023covariate for related discussions. Recent work has developed such methods of covariate adjustment for specific parameters of interest (see, for instance, bai2024covariate; bai2024inference-1; bai2025inference; cytrynbaum2023covariate), but we leave the development of a method of covariate adjustment which applies at the level of generality considered in this paper to future work.
@@ Appendix A Proofs of Main Results
By the proof of Theorem 5.21 in van_der_vaart1998asymptotic, to show (12), it suffices to show
@@
To accomplish this, we study $\mathbb{L}_{n}^{(s)}(\theta)$ for $1\leq s\leq d_{\theta}$ separately. It follows from Assumption 3.3(c)–(d), Proposition 8.11 in kosorok2008introduction, and the arguments to establish (48) that
@@
Therefore, since $\hat{\theta}_{n}\stackrel{{\scriptstyle P}}{{\to}}\theta_{0}$ by Lemma A.8, to show (25) it suffices to argue that for every $\epsilon>0$ and every sequence $\delta_{n}\downarrow 0$ (p.89 of van_der_vaart1996weak),
@@
Fix any sequence $\tilde{\delta}_{n}\downarrow 0$. For every $n$, there exists $n^{\prime}$ such that $\{\theta\in\Theta^{\ast}:\|\theta-\theta_{0}\|<\delta_{n^{\prime}}\}\subseteq\{\theta\in\Theta^{\ast}:\rho_{Q}(\theta,\theta_{0})<\tilde{\delta}_{n}\}$. By Proposition C.1 in han2021complex,
@@
where the convergence follows from Assumption 3.3(e) and Corollary 2.3.12 in van_der_vaart1996weak. We then obtain (26) by Markov’s inequality.
@@
Recall $\Omega^{(s)}(X_{i})$ is the $s$th component of $\Omega(X_{i})$. Next, it follows from Assumption 3.1, 3.3(f), and equation (12.3) in lehmann2022testing that for $1\leq s\leq d_{\theta}$,
@@
In what follows, we will show (27). The proof of (28) will follow from similar steps, but with the calculations below replaced by the ones in the proof of Lemmas C.2–C.3 in bai2024inference, along with Assumptions 3.43.5. To that end, we show for $1\leq s\leq d_{\theta}$,
@@
Recall that $P_{n}$ denotes the distribution of the observed data $(X^{(n)},A^{(n)},R^{(n)})$, and $Q$ denotes the marginal distribution of the vector $(R_{i}(1),R_{i}(0),X_{i})$. Note that any treatment assignment mechanism $A^{(n)}$ satisfying Assumption 2.1 can be represented as a function of $X^{(n)}$ and some additional exogenous randomization device $U_{n}\in\mathbb{R}$. Let $p_{n}^{U_{n}}$ denote the density function for $U_{n}$ with respect to a dominating measure $\mu^{U}$. In what follows, we consider a family $\{Q_{t}:t\in\mathbf{R}^{d_{\theta}}\}$ of marginal distributions indexed by $t$, and let $q_{t}^{X}$ denote the density function for $X_{i}$ with respect to a dominating measure $\mu^{X}$, $q_{t}^{R(a)|X}(r|x)$ denote the conditional density of $R_{i}(a)$ given $X_{i}$ with respect to a dominating measure $\mu^{R}$. With some abuse of notation, continue letting $P_{t,n}$ denote the distribution of $(U_{n},X^{(n)},R^{(n)})$. We require that $Q_{0}=Q$ and $P_{0,n}=P_{n}$ and define $q^{X}=q_{0}^{X}$ and $q^{R(a)|X}=q_{0}^{R(a)|X}$. As a consequence, the density function of $P_{t,n}$ is given by
@@
(a) and (b) follow from Lemma 14.2.1 in lehmann2022testing. (c) follows from the same lemma. In order to show (d), fix $t_{n}\to 0$. Note (34) and Markov’s inequality imply that along a subsequence $t_{n_{k}}$,
@@
for $Q$-almost every $x$. Along that subsequence, another application of Lemma 14.2.1 in lehmann2022testing implies (d).
@@
We use a vector version of the construction in in Example 25.16 in van_der_vaart1998asymptotic. Let $k(x)$ be any strictly positive function that is bounded from above and away from zero with a bounded derivative such that $k(0)=k^{\prime}(0)=1$; for example, take $k(x)=2(1+e^{-2x})^{-1}$. Define
@@
The quadratic mean differentiability requirement in (33) follows from Lemma 7.6 in van_der_vaart1998asymptotic. Next, for each $x\in\mathbf{R}^{d_{x}}$, we define
@@
To apply Vitali’s theorem (Proposition 2.29 in van_der_vaart1998asymptotic), it suffices to show that
@@
The first result follows from Theorem 3.1 of armstrong2022asymptotic. The second result follows from Lemma A.6 given Assumption 4.1 and the assumption that each component of $I^{R(a)|X}(x)$ is integrable, noting that $E[s_{g}(X,1,R(1))-s_{g}(X,0,R(0))|X]=0$.
@@
Note that $t^{\prime}s_{g}$ appears in the expansion of the log-likelihood ratio between $P_{t/\sqrt{n},g}$ and $P$. To align our setting with Theorem 3.11.2 in van_der_vaart1996weak, we first characterize the adjoint map (viewed as a mapping into $\mathcal{M}_{g}$) of the function $v\mapsto E[\psi^{*}v]\in\mathbb{R}^{d_{\theta}}$, where $v\in\mathcal{M}_{g}$. To that end, implicitly identifying each $b\in\mathbf{R}^{d_{\theta}}$ with the functional $b^{*}:\mathbb{R}^{d_{\theta}}\rightarrow\mathbb{R}$ given by $x\mapsto b^{\prime}x$, we construct a $w(b)\in\mathbf{R}^{d_{\theta}}$ such that
@@
Here we use the assumption that $I$ is nonsingular. It then follows from the local asymptotic normality established in Lemma A.3 and Theorem 3.11.2 in van_der_vaart1996weak that
@@
The formula for the derivative therefore follows from the implicit function theorem (in particular, because we have assumed the existence of $\theta(P_{\nu})$ along the path, it follows from the last part of the proof of Theorem 3.2.1 in krantz2013implicit). The second equality follows from Lemma A.7 together with Condition A.1.
@@
This assumption guarantees that every element satisfying Condition A.1 has a corresponding path for which we can apply Lemma A.5. A similar assumption appears in chen2018overidentification (see their Assumption 4.1(iv)). Note that a simple sufficient condition for the first part of Assumption A.1 is that $m(x,a,r,\theta(P))$ is a bounded function in $(x,r)$ on the support of $(X,R(a))$. The second part of Assumption A.1 can be verified easily in specific examples (see, for instance, Examples 2.12.5 in the main text). Alternatively, Assumption A.1 could be avoided by assuming that we can differentiate under the integral in the final step of the proof of Lemma A.5, from which we would immediately obtain the expression for the pathwise derivative. See, for instance, newey1994asymptotic and chen2008semiparametric.
@@
As a result, one can verify the Lindeberg condition conditional on $X^{(n)}$ and $A^{(n)}$ as in the proof of Lemma S.1.4 of bai2022inference, and obtain
$\rho(\mathbb{C}_{3,n},N(0,V_{3,f})|X^{(n)},A^{(n)})\xrightarrow{P}0~.$
The proof can then be completed by the subsequencing argument at the end of the proof of Lemma S.1.4 of bai2022inference.
@@ A.4 Auxiliary Lemmas
It follows from Assumption 3.3(a) and Theorem 5.9 in van_der_vaart1998asymptotic that we only need to establish for each $1\leq s\leq d_{\theta}$,
@@
To begin, note it follows from Assumption 3.3(d) and the dominated convergence theorem that if $m_{s}(x,a,r,\theta_{m})\allowbreak\to m_{s}(x,a,r,\theta)$ as $m\to\infty$ for $\{\theta_{m}\}\subset\Theta^{\ast}$, then $E_{P}[m_{s}(X_{i},A_{i},R_{i},\allowbreak\theta_{m})]\to E_{P}[m_{s}(X_{i},A_{i},R_{i},\theta)]$. Here, the dominating function exists by Problem 2.4.1 in van_der_vaart1996weak. Assumption 3.3(c) then implies
@@
where the first equality follows from (48), the first inequality follows from the triangle inequality, the second equality follows because $\sum_{1\leq i\leq n}A_{i}=n\eta$ and $\sum_{1\leq i\leq n}(1-A_{i})=n(1-\eta)$, and the last inequality follows from Proposition C.1 in han2021complex. The convergence follows from Assumption 3.3(e) and an application of the backward submartingale convergence theorem (see, for instance, Theorem 12.30 in le2022measure), as detailed in the proof of Theorem 3.1 in han2021complex. The desired result in (47) then follows by Markov’s inequality.
@@
The desired result in (51) follows from similar arguments to those in the proof of equation (S.27) in bai2022inference, where the uniform integrability condition is replaced by Assumption 3.5(a). To show (52), note

### v8

@@ 6 Recommendations for Empirical Practice
We conclude with some recommendations for empirical practice based on our theoretical results. Overall, our findings highlight the general benefit of highly stratified designs for designing efficient experiments: highly stratified experiments “automatically” perform fully-efficient covariate adjustment for a large class of interesting parameters. This finding generalizes similar observations made by Bai et al. (2022), Bai (2022) and Cytrynbaum (2023b) for the special case of estimating the ATE.
Our simulation evidence suggests, however, that highly stratified experiments may produce less precise estimates than (correctly specified) covariate adjustment when the the dimension of $X_{i}$ is large relative to the sample size. For this reason, we recommend that practitioners construct their blocks using a subset of the baseline covariates that they believe have the highest explanatory power in terms of the nonparametric $R^{2}$ in (19); the pre-treatment measure of the outcomes of interest, for example, is typically believed to be one such covariate (see, in particular, Bruhn and McKenzie, 2009). The experimental data can then be analyzed efficiently using an unadjusted method-of-moments estimator.
If one wishes to perform covariate adjustment with additional covariates beyond those used for blocking, then this can be done ex-post. As discussed in Remark 4.2, the scope for improvement from covariate adjustment is limited by the nonparametric $R^{2}$ from the regression of the moment functions on the additional covariates, conditional on the ones used for matching; if one has already matched on the covariates with the highest explanatory power, then the potential gain in efficiency from adjusting for these additional covariates may be limited. We further caution that care must be taken to ensure that the adjustment is performed in such a way that it guarantees a gain in efficiency: see Bai et al. (2024a) and Cytrynbaum (2023a) for related discussion. Recent work has developed such methods of covariate adjustment for specific parameters of interest (see, for instance, Bai et al., 2024a, b, 2025a; Cytrynbaum, 2023a), but we leave the development of a method of covariate adjustment which applies at the level of generality considered in this paper to future work.
@@ Appendix A Proofs of Main Results
By the proof of Theorem 5.21 in van der Vaart (1998), to show (12), it suffices to show
@@
To accomplish this, we study $\mathbb{L}_{n}^{(s)}(\theta)$ for $1\leq s\leq d_{\theta}$ separately. It follows from Assumption 3.3(c)–(d), Proposition 8.11 in Kosorok (2008), and the arguments to establish (48) that
@@
Therefore, since $\hat{\theta}_{n}\stackrel{{\scriptstyle P}}{{\to}}\theta_{0}$ by Lemma A.8, to show (25) it suffices to argue that for every $\epsilon>0$ and every sequence $\delta_{n}\downarrow 0$ (p.89 of van der Vaart and Wellner, 1996),
@@
Fix any sequence $\tilde{\delta}_{n}\downarrow 0$. For every $n$, there exists $n^{\prime}$ such that $\{\theta\in\Theta^{\ast}:\|\theta-\theta_{0}\|<\delta_{n^{\prime}}\}\subseteq\{\theta\in\Theta^{\ast}:\rho_{Q}(\theta,\theta_{0})<\tilde{\delta}_{n}\}$. By Proposition C.1 in Han and Wellner (2021),
@@
where the convergence follows from Assumption 3.3(e) and Corollary 2.3.12 in van der Vaart and Wellner (1996). We then obtain (26) by Markov’s inequality.
@@
Recall $\Omega^{(s)}(X_{i})$ is the $s$th component of $\Omega(X_{i})$. Next, it follows from Assumption 3.1, 3.3(f), and equation (12.3) in Lehmann and Romano (2022) that for $1\leq s\leq d_{\theta}$,
@@
In what follows, we will show (27). The proof of (28) will follow from similar steps, but with the calculations below replaced by the ones in the proof of Lemmas C.2–C.3 in Bai et al. (2024c), along with Assumptions 3.43.5. To that end, we show for $1\leq s\leq d_{\theta}$,
@@
Recall that $P_{n}$ denotes the distribution of the observed data $(X^{(n)},A^{(n)},R^{(n)})$, and $Q$ denotes the marginal distribution of the vector $(R_{i}(1),R_{i}(0),X_{i})$. Note that any treatment assignment mechanism $A^{(n)}$ satisfying Assumption 2.1 can be represented as a function of $X^{(n)}$ and some additional exogenous randomization device $U_{n}\in\mathbf{R}$. Let $p_{n}^{U_{n}}$ denote the density function for $U_{n}$ with respect to a dominating measure $\mu^{U}$. In what follows, we consider a family $\{Q_{t}:t\in\mathbf{R}^{d_{\theta}}\}$ of marginal distributions indexed by $t$, and let $q_{t}^{X}$ denote the density function for $X_{i}$ with respect to a dominating measure $\mu^{X}$, $q_{t}^{R(a)|X}(r|x)$ denote the conditional density of $R_{i}(a)$ given $X_{i}$ with respect to a dominating measure $\mu^{R}$. With some abuse of notation, continue letting $P_{t,n}$ denote the distribution of $(U_{n},X^{(n)},R^{(n)})$. We require that $Q_{0}=Q$ and $P_{0,n}=P_{n}$ and define $q^{X}=q_{0}^{X}$ and $q^{R(a)|X}=q_{0}^{R(a)|X}$. As a consequence, the density function of $P_{t,n}$ is given by
@@
(a) and (b) follow from Lemma 14.2.1 in Lehmann and Romano (2022). (c) follows from the same lemma. In order to show (d), fix $t_{n}\to 0$. Note (34) and Markov’s inequality imply that along a subsequence $t_{n_{k}}$,
@@
for $Q$-almost every $x$. Along that subsequence, another application of Lemma 14.2.1 in Lehmann and Romano (2022) implies (d).
@@
We use a vector version of the construction in in Example 25.16 in van der Vaart (1998). Let $k(x)$ be any strictly positive function that is bounded from above and away from zero with a bounded derivative such that $k(0)=k^{\prime}(0)=1$; for example, take $k(x)=2(1+e^{-2x})^{-1}$. Define
@@
The quadratic mean differentiability requirement in (33) follows from Lemma 7.6 in van der Vaart (1998). Next, for each $x\in\mathbf{R}^{d_{x}}$, we define
@@
To apply Vitali’s theorem (Proposition 2.29 in van der Vaart, 1998), it suffices to show that
@@
The first result follows from Theorem 3.1 of Armstrong (2022). The second result follows from Lemma A.6 given Assumption 4.1 and the assumption that each component of $I^{R(a)|X}(x)$ is integrable, noting that $E[s_{g}(X,1,R(1))-s_{g}(X,0,R(0))|X]=0$.
@@
Note that $t^{\prime}s_{g}$ appears in the expansion of the log-likelihood ratio between $P_{t/\sqrt{n},g}$ and $P$. To align our setting with Theorem 3.11.2 in van der Vaart and Wellner (1996), we first characterize the adjoint map (viewed as a mapping into $\mathcal{M}_{g}$) of the function $v\mapsto E[\psi^{*}v]\in\mathbf{R}^{d_{\theta}}$, where $v\in\mathcal{M}_{g}$. To that end, implicitly identifying each $b\in\mathbf{R}^{d_{\theta}}$ with the functional $b^{*}:\mathbf{R}^{d_{\theta}}\rightarrow\mathbf{R}$ given by $x\mapsto b^{\prime}x$, we construct a $w(b)\in\mathbf{R}^{d_{\theta}}$ such that
@@
Here we use the assumption that $I$ is nonsingular. It then follows from the local asymptotic normality established in Lemma A.3 and Theorem 3.11.2 in van der Vaart and Wellner (1996) that
@@
The formula for the derivative therefore follows from the implicit function theorem (in particular, because we have assumed the existence of $\theta(P_{\nu})$ along the path, it follows from the last part of the proof of Theorem 3.2.1 in Krantz and Parks (2013)). The second equality follows from Lemma A.7 together with Condition A.1.
@@
This assumption guarantees that every element satisfying Condition A.1 has a corresponding path for which we can apply Lemma A.5. A similar assumption appears in Chen and Santos (2018) (see their Assumption 4.1(iv)). Note that a simple sufficient condition for the first part of Assumption A.1 is that $m(x,a,r,\theta(P))$ is a bounded function in $(x,r)$ on the support of $(X,R(a))$. The second part of Assumption A.1 can be verified easily in specific examples (see, for instance, Examples 2.12.5 in the main text). Alternatively, Assumption A.1 could be avoided by assuming that we can differentiate under the integral in the final step of the proof of Lemma A.5, from which we would immediately obtain the expression for the pathwise derivative. See, for instance, Newey (1994) and Chen et al. (2008).
@@
As a result, one can verify the Lindeberg condition conditional on $X^{(n)}$ and $A^{(n)}$ as in the proof of Lemma S.1.4 of Bai et al. (2022), and obtain
$\rho(\mathcal{L}(\mathbb{C}_{3,n}|X^{(n)},A^{(n)}),N(0,V_{3,f}))\xrightarrow{P}0~.$
The proof can then be completed by the subsequencing argument at the end of the proof of Lemma S.1.4 of Bai et al. (2022).
@@ A.4 Auxiliary Lemmas
It follows from Assumption 3.3(a) and Theorem 5.9 in van der Vaart (1998) that we only need to establish for each $1\leq s\leq d_{\theta}$,
@@
To begin, note it follows from Assumption 3.3(d) and the dominated convergence theorem that if $m_{s}(x,a,r,\theta_{m})\allowbreak\to m_{s}(x,a,r,\theta)$ as $m\to\infty$ for $\{\theta_{m}\}\subset\Theta^{\ast}$, then $E_{P}[m_{s}(X_{i},A_{i},R_{i},\allowbreak\theta_{m})]\to E_{P}[m_{s}(X_{i},A_{i},R_{i},\theta)]$. Here, the dominating function exists by Problem 2.4.1 in van der Vaart and Wellner (1996). Assumption 3.3(c) then implies
@@
where the first equality follows from (48), the first inequality follows from the triangle inequality, the second equality follows because $\sum_{1\leq i\leq n}A_{i}=n\eta$ and $\sum_{1\leq i\leq n}(1-A_{i})=n(1-\eta)$, and the last inequality follows from Proposition C.1 in Han and Wellner (2021). The convergence follows from Assumption 3.3(e) and an application of the backward submartingale convergence theorem (see, for instance, Theorem 12.30 in Le Gall, 2022), as detailed in the proof of Theorem 3.1 in Han and Wellner (2021). The desired result in (47) then follows by Markov’s inequality.
@@
The desired result in (51) follows from similar arguments to those in the proof of equation (S.27) in Bai et al. (2022), where the uniform integrability condition is replaced by Assumption 3.5(a). To show (52), note
@@
Appendix B Details for Simulations
B.1 Local Average Treatment Effect
[…]
We consider the following three estimators for the LATE:
  • Unadjusted Estimator:
  • Adjusted Estimator 1:
Similarly to Section 5.1, $\hat{\theta}_{n}^{\rm unadj}$ solves (6) for the moment condition given in (5). The second and third estimators are covariate adjusted estimators which can be obtained as two-step method of moments estimators from solving an “augmented” version of the moment condition (5) (see, for instance, Chernozhukov et al., 2018; Jiang et al., 2022a).
B.2 Additional Simulation Tables
Table 3 reports the coverage and average length of confidence intervals constructed using the estimator described in Section 3.1, again computed across $4000$ Monte Carlo replications. We find that the confidence intervals have appropriate coverage at all sample sizes and for any number of covariates. The average length of the intervals follows a pattern similar to what was observed for the MSE, again because the first 4 covariates are much stronger predictors of the control outcome than the last 4 covariates, which are almost uninformative. Although we found in Table 1 that the MSE of the matched tuples design was worse than that of matched pairs with 8 covariates, this does not seem to translate to longer confidence intervals for matched tuples: this may be due to the fact that the matched tuples variance estimator is a “within block” estimator, as described in Section 3.1. Similar findings have been reported by Bai et al. (2024c) for related problems in factorial designs.
i.i.d. assignment Matched pairs Matched tuples
$n=100$
ATE 2 0.0083 0.0065 0.0051 0.0212 0.0068
4 0.0083 0.0090 0.0090 0.0073 0.0124
8 0.0083 0.0134 0.0131 0.0010 0.0209
LATE 2 0.0905 0.0436 0.0396 0.0341 0.0076
4 0.0905 0.0276 0.0264 0.0301 0.0427
8 0.0905 0.0401 0.0376 0.0264 0.0297
$n=200$
ATE 2 0.0075 0.0068 0.0074 0.0084 0.0081
4 0.0075 0.0052 0.0067 0.0040 0.0040
8 0.0075 0.0045 0.0055 0.0145 0.0051
LATE 2 0.0545 0.0326 0.0339 0.0361 0.0328
4 0.0545 0.0182 0.0217 0.0123 0.0144
8 0.0545 0.0159 0.0180 0.0447 0.0216
$n=400$
ATE 2 0.0001 0.0011 0.0009 0.0041 0.0055
4 0.0001 0.0001 0.0001 0.0001 0.0016
8 0.0001 0.0006 0.0006 0.0056 0.0041
LATE 2 0.0142 0.0064 0.0059 0.0179 0.0198
4 0.0142 0.0012 0.0015 0.0004 0.0020
8 0.0142 0.0021 0.0019 0.0129 0.0064
$n=1000$
ATE 2 0.0021 0.0037 0.0039 0.0014 0.0025
4 0.0021 0.0021 0.0022 0.0020 0.0006
8 0.0021 0.0021 0.0023 0.0023 0.0027
LATE 2 0.0118 0.0120 0.0123 0.0019 0.0083
4 0.0118 0.0055 0.0057 0.0051 0.0007
8 0.0118 0.0054 0.0058 0.0057 0.0059
$n=2000$
ATE 2 0.0013 0.0010 0.0010 0.0011 0.0025
4 0.0013 0.0015 0.0015 0.0003 0.0021
8 0.0013 0.0015 0.0015 0.0001 0.0014
LATE 2 0.0062 0.0035 0.0035 0.0018 0.0050
4 0.0062 0.0037 0.0038 0.0011 0.0055
8 0.0062 0.0038 0.0037 0.0008 0.0032
Matched pairs Matched tuples
# of covariates ($T$) Coverage CI length Coverage CI length
$n=100$
ATE 2 0.9403 2.3191 0.9490 2.3224
4 0.9465 2.2744 0.9485 2.2830
8 0.9487 2.3532 0.9395 2.3672
LATE 2 0.9590 6.4075 0.9590 6.3865
4 0.9620 6.3184 0.9615 6.3176
8 0.9610 6.5497 0.9527 6.5311
$n=200$
ATE 2 0.9495 1.6401 0.9497 1.6418
4 0.9485 1.5990 0.9440 1.6049
8 0.9503 1.6529 0.9425 1.6617
LATE 2 0.9550 4.4046 0.9563 4.4264
4 0.9557 4.3315 0.9530 4.3381
8 0.9597 4.4833 0.9503 4.5038
$n=400$
ATE 2 0.9487 1.1602 0.9483 1.1614
4 0.9497 1.1255 0.9455 1.1285
8 0.9495 1.1609 0.9445 1.1667
LATE 2 0.9530 3.0955 0.9493 3.0963
4 0.9547 3.0122 0.9517 3.0110
8 0.9555 3.1284 0.9505 3.1239
$n=1000$
ATE 2 0.9485 0.7349 0.9550 0.7352
4 0.9500 0.7086 0.9505 0.7102
8 0.9417 0.7285 0.9430 0.7318
LATE 2 0.9507 1.9468 0.9560 1.9452
4 0.9515 1.8836 0.9507 1.8857
8 0.9467 1.9469 0.9463 1.9480
$n=2000$
ATE 2 0.9490 0.5194 0.9530 0.5193
4 0.9483 0.4998 0.9485 0.5004
8 0.9505 0.5121 0.9483 0.5139
LATE 2 0.9495 1.3714 0.9530 1.3721
4 0.9493 1.3250 0.9483 1.3258
8 0.9533 1.3663 0.9505 1.3658